how much heat is given out when 50g of steam at 1000c cool to water at 280c? (specific latent heat of vaporization of water =2.3x106jkg-1)
1
Expert's answer
2020-06-05T11:58:06-0400

Heat is given out "Q = m\times C_{p}\times \Delta t"


"Q = 0.05 \times 2.3\times10^{-3}\times (1000-280) =82800 kJ"